OpenStudy (anonymous):

(f+g)(x)=(x+9)+(x-9) please explain steps

OpenStudy (anonymous):

add \[x+9+x-9=2x\] i think what was the question?

OpenStudy (anonymous):

find (fx+g)(x)

OpenStudy (anonymous):

i got x^2-81

OpenStudy (anonymous):

that would be \(f\times g(x)\)

OpenStudy (anonymous):

add means add them

OpenStudy (anonymous):

the answer is 2x and i dont understand y

OpenStudy (anonymous):

do you have: f(x) = and g(x) = ??

OpenStudy (anonymous):

no please explain

OpenStudy (anonymous):

the problem you posted askes for the sum of f(x) + g(x). I was wondering if they gave you what each of these functions are...

OpenStudy (anonymous):

no thats all that was given

OpenStudy (anonymous):

oh nvm this stupid thing confused me with the parantheses i thought i was supposed to foil because of the aprantheses which confuse me wow........ such a fail i get it now

OpenStudy (savvy):

(f+g)x means f(x) + g(x) however you have been directly given (f+g)x = (x+9)+(x-9) open the brackets and write x + 9 + x - 9 rearrange x + x + 9 - 9 cmbine like terms. (x + x) + (9-9) 9-9 = 0 and x +x= 2x so final answer is 2x + 0 = 2x